# Automatic Combination of Sample Selection Strategies for Few-Shot Learning

### 05 Feb 2024 v1

@@ 4.1 Impact of Single-Property Selection Strategies
In this section, our goal is to answer the following research question: RQ1: What is the impact of single-property selection strategies on the few-shot learning? The objective is to analyse the performance increase of the different selection strategies across various datasets and few-shot learning approaches. The results are shown in Figure 1 and Table 2, aggregated over all the datasets (the detailed de-aggregated results are included in Appendix F). We present the results separately for image and text datasets, as well as for the gradient few-shot learning and in-context learning, as we observed a strong dependence of the strategy benefit on the modality and few-shot learning approach used.
Majority of selection strategies fail to outperform the Classic selection in case of gradient few-shot learning. In most cases, the strategies lead to performance on par or even lower than the Classic selection (for example hard samples on image datasets or diverse samples on text datasets). Only specific strategies consistently lead to better performance (e.g., easy to learn samples on image datasets, or Glister on text datasets). As such, the random sample selection represents a strong baseline for the gradient few-shot learning.
Sample selection is more beneficial for in-context learning. Larger number of strategies leads to increase in performance, while only few strategies, such as the popular selection based on similarity or diversity, under-perform the Classic selection. At the same time, the increase in performance is higher than with gradient few-shot learning, although with a larger variance in results. In contrast to gradient few-shot learning, the hard to learn samples provide the most benefit for in-context learning. The reason may be that hard to learn samples provide a lot of information beneficial for the classification, which cannot be easily leveraged in the gradient few-shot learning due to the samples being hard to learn. However, as there is no explicit “learning” in the in-context learning use of large language models, such samples can be more easily leveraged.
Learnability property is a strong indicator of quality for few-shot learning. The strategies that choose based on how easy it is to learn the samples (Cartography) or how often they are forgotten (Forgetting) consistently lead to overall highest increase in performance. The observed improvement is on average as high as percentage point for gradient few-shot learning and percentage points for in-context learning.
4.2 ACSESS Results and Characteristics
In this section, we answer the following research question: RQ2: How does the ACSESS method perform in comparison to the single-property strategies? Besides analysing the benefit and comparing with other strategies, we analyse the influence of the identified strategies and of the different weighting schemes.
The ACSESS method consistently leads to performance increase across all few-shot learning approaches. Our proposed method outperforms all the investigated single-property strategies, leading to an average performance increase of up to percentage points for gradient few-shot learning and percentage points for in-context learning. Comparing with the in-context specific strategy LENS (Li & Qiu, 2023) (see Figure 2 for comparison), which also outperforms all single-property strategies, we observe the increase of ACSESS is statistically more significant (p-value of using the Wilcoxon signed-rank test).
The ACSESS method identifies samples with complementary properties, with higher priority on learnability. The identified selection strategies that are used in the combination are different for the different data modalities (image vs. text) and the type of few-shot learning (gradient vs. in-context learning). For the image datasets, the combination of easy to learn samples (Cartography) that are located on the decision boundary (Margin) provides the most benefit. For the gradient few-shot learning approaches on text datasets, the combination of easy and ambiguous samples (Cartography) that are least often forgotten (Forgetting) on the decision boundary (Margin) that provide additional informativeness from the Graph-Cut strategy provides the most benefit. In case of in-context learning, the combination of hard to learn samples (Cartography) that are least often forgotten (Forgetting) on the decision boundary (Margin) and with highest Entropy provide the most benefit. Further analysing the identified strategies and their weights, we find the ACSESS method gives higher importance to samples that are learnable, lower importance to informative samples (but with multiple sources for both), while the strategies that focus on representativeness are not included at all.
The dataset and approach specific weighting leads to the best performance. The only exception is the Few-Shot Fine-Tuning approach, where the uniform weighting shows the most benefit. At the same time, the weighting strategy that incorporates the random selection often leads to worse performance and higher variance in the results. Even though the uniform combination is computationally less expensive than the weighted combination, its performance increase is only slightly lower (average difference of ). As such, the uniform weighting represents a good trade-off between the performance increase and the computation costs, making it in general a better choice.
4.3 Effects of Variable Number of Shots
In this section, we answer the following research question: RQ3: How does the benefit of sample selection change based on the number of shots? Our aim is to investigate how the benefit of sample selection strategies changes as we increase or decrease the number of shots that we select. In addition, we investigate whether the few-shot learning approaches themselves benefit from larger number of shots. The results of this evaluation are included in Figure 3, comparing the benefit of the ACSESS strategy and Random selection over the main experimental setup (5-shot setting).
The selection strategies provide more benefit on lower number of shots. On the 1-shot setting, the difference in accuracy between the Random strategy and the ACSESS strategy can be as high as , or percentage points respectively for the few-shot learning on image data and text data, and in-context learning on text data. Increasing the number of selected shots leads to diminished impact, as the ACSESS strategy regresses to random selection when selecting more than 30-40 shots. For more detailed analysis we compare all the investigated strategies in the 5-shot and 10-shot setting (distribution of results is shown in Figure 5 in Appendix E). This comparison further reinforces our findings, as we find that the difference to random selection is statistically significantly higher in the 5-shot (p-value of using the Wilcoxon signed-rank test).
Impact of selection strategies and further increase of the number of shots is negligible after certain point. For the gradient few-shot learning, the performance after approximately 50 shots remains more or less constant. For the in-context learning, the performance even decreases significantly after 20-25 shots, achieving similar performance to the 1-shot setting when using 50-shots. The main reason for this drop-off may be the limited context size of the large language models, making the model degenerate to random guessing as the cross-over is text length dependent. As such, the focus on selecting high quality samples is paramount when only few labelled samples are available (e.g., working with limited labelling budget), while the random selection can be safely used with enough labels. Furthermore, labelling more samples and using them as additional shots for few-shot learning may not lead to better performance.
4.4 Additional Insights to Sample Selection Strategies
@@
In this section, we provide additional insights from ablation studies. We are mainly interested how the factors such as dataset and model affect the sample selection (based on dataset specific results from Appendix F) and how sensitive the strategies are to their random components (based on the variance from repeated strategy runs in Figure 4).
@@
Specific strategies show sensitivity to the effects of random seed. Changing the random seed for specific strategies may often lead to different set of selected samples and different performance impact. The only exception are some of the core set selection strategies (such as Glister or the Graph-Cut), which consistently lead to the selection of the same set of samples. On the other hand, strategies with random initial condition, such as the similarity selection or some of the active learning strategies, are more sensitive to the effects of the random seed. The ACSESS method shows lower sensitivity to repeated runs (with the exception of weighting with random element) but still may lead to different sets of samples based on the random seed.
@@
Strong dataset dependence for selection strategies. The strategy that performs well on one dataset (such as DeepFool on the HUM_ACT.ACT_410 dataset) may perform quite poorly on other datasets (such as MCR.PRT or MCR.PNU). Such strategies cannot be used on any dataset without further evaluation. Only few strategies perform consistently on the majority of the datasets, such as Herding showing poor performance on all image datasets, or Cartography showing good performance across all datasets and models. We do not observe a significant dataset dependence of the ACSESS method, as it identifies the well-performing strategies and leverages their strengths for the specific dataset.
@@
Strong model dependence and sensitivity to selected samples. The strategy performing well for one model (GraNd on MCR.PNU for MAML, or KCenter on 20 News for Zephyr) may result in a decrease of performance for other models (GraNd for Few-Shot Fine-Tuning, or KCenter for Mistral on the same setting). Overall, the selected samples resulting from different strategy or from a repeated run of the same strategy affect the Few-Shot Fine-Tuning less, leading to smaller benefit. At the same time, the MAML method and the in-context learning models are more affected by the selected samples, leading to more varied results and often higher performance increase. Such sensitivity to sample selection was observed in previous works as well (Pecher et al., 2023; Köksal et al., 2022; Zhang et al., 2022; Li & Qiu, 2023). Curiously, Prototypical Networks gain quite large performance increase from sample selection, but show low sensitivity to the selected samples and variance from repeated runs of sample selection. As such, sample selection is more important for MAML and In-Context Learning and less important for Few-Shot Fine-Tuning.
5 Conclusion
In this paper, we investigate the impact of larger number of different sample selection strategies on the success of few-shot learning. To leverage strengths of these strategies, we propose the ACSESS method for automatic combination of the sample selection strategies to select samples with complementary properties. The experimental results show that although the individual sample selection strategies may not always be beneficial, their combination using our proposed method consistently leads to better performance. Curiously, we observe that the learnability of the samples is more important for few-shot learning than their informativeness or representativeness. Finally, analysing the behaviour of the sample selection in regards to the number of selected shots, we find that the sample selection provides the most benefit on low number of shots, while at the higher number of shots all strategies regress to random guessing. In addition, after certain point, increasing the number of available samples no longer leads to improvement in performance of few-shot learning, but may lead to decrease in performance in some cases, especially for in-context learning.
Acknowledgements
This work was partially supported by the projects funded by the European Union under the EU Horizon 2020: TAILOR, GA No. 952215, by the European Union under the Horizon Europe: vera.ai, GA No. 101070093, AI-CODE, GA No. 101135437; and by Modermed, a project funded by the Slovak Research and Development Agency, GA No. APVV-22-0414.
Part of the research results was obtained using the computational resources procured in the national project National competence centre for high performance computing (project code: 311070AKF2) funded by European Regional Development Fund, EU Structural Funds Informatization of Society, Operational Program Integrated Infrastructure.
Ethical Considerations and Impact Statement
The experiments in this paper work with publicly available datasets, citing the original authors of each dataset. All the datasets are used in accordance to their license and terms and conditions of use. We do not work with any personally identifiable information or offensive content and perform no crowdsourcing or further data annotation. In addition, we are not aware of any potential ethical harms or negative societal impacts of our work, apart from the ones related to the advancement of the field of Machine Learning and Few-Shot learning. Finally, we follow the license terms for all the models we use – all models and datasets allow their use as part of research. It is possible the large language models (Mistral and Zephyr) we used contain biases and may generate potentially offensive or harmful content. However, the authors of the models already reduce this potential bias as much as possible when training the models, while at the same time we limit the output to few tokens and do not release any output of these models, which should further reduce the potential bias and negative impact.
Impact Statement: CO2 Emissions Related to Experiments
The experiments presented in this paper used higher compute resources in order to evaluate all the sample selection strategies for all the different few-shot learning approaches. Overall, the experiments were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCOeq/kWh. A cumulative of 2000 hours of computation was performed on hardware of type A100 PCIe 40GB (TDP of 250W). Total emissions are estimated to be 216 kgCOeq of which 0 percent were directly offset. These estimations were conducted using the MachineLearning Impact calculator presented in (Lacoste et al., 2019). Whenever possible, we tried to reduce the compute resources used as much as possible. The most compute resources were used by the large language models – Mistral-7B and Zephyr-7B and their further tuning. To reduce the computation costs, we use the 4-bit quantised versions of these models and in cases when the further tuning is required, we fine-tune only the last classification layer (keeping the rest frozen). In addition, for gradient few-shot learning approaches, we opt for smaller models that do not require such an extensive computation costs, while still providing relevant results. Finally, whenever possible, we reuse the already performed computations for the selection strategies, such as generating the feature representation for the different samples only a single time and re-using it whenever possible (e.g., using pre-trained models without further tuning for these representations).
Appendix A Limitations
In the experiments for evaluating the impact of different sample selection strategies, we focused on such settings, where the impact of sample selection may be reduced — despite that we observed significant differences in the achieved performance. First of all, in within-domain few-shot learning, the samples are selected from the same domain, thus the difference in impact and role of the individual samples may not be as high. We may expect that the impact of sample selection in cross-domain few-shot learning may be higher and lead to even larger performance deviations. In addition, we focus on a balanced-shot setting, where we select the same number of shots for each class, leading to lower impact of strategies for different classes. These additional settings are out of scope of this paper and may be a subject of future research.
To diminish undesirable randomness effects, we report the results as an average over 5 repeats of data splitting, 10 repeated runs of the selection strategies (resulting in 50 overall evaluation runs), with each run evaluated on 600 tasks. Such setting limits the variance in results, but may also obscure the benefit of selection strategies to a certain extent. On the other hand, although we focus only on a 5-way few-shot learning setting (using only 5 classes per task), we do not expect significant change in result for most of the few-shot learning approaches when using higher number of classes (such as 10-way or 20-way classification). The only exception is the in-context learning, where we observed the dependence on input size and context width of the models. With higher number of classes, the observed performance of large language models would start to deteriorate to the performance that can be achieved with a low number of shots.
Similarly to other research papers focusing on sample selection and few-shot learning via in-context learning, we assume an availability of large-enough annotated dataset from which we select the sample using different strategies. This also includes the use of larger validation or held-out set when selecting strategies. If we considered the true few-shot learning setting, only the active learning strategies (i.e., Entropy, Margin and Least Confidence) could be used, as they work with unlabelled data only, or a different approach for sample selection should be employed, such as using weak-labels to select the best samples, or assign every single class to all the labels and perform the sample selection that way to select both best performing samples as well as the class for them (as it is done in (Chang & Jia, 2023)).
Before the experiments, we run hyperparameter optimisation for the gradient few-shot learning models and perform basic prompt-engineering for the in-context learning approaches. We follow the recommendations for good prompt engineering, using the meta-tags defined for each model (which lead to better performance of the models), while also not using overly-simple or overly-complicated prompts. The prompts we use are created based on dataset description, the prompts used in related works and the formats recommended for different models (e.g., taking inspiration from (Sun et al., 2023)). As we do not perform any extensive or automatic prompt-engineering, these prompts may lead to lowered performance and in some cases also lower impact of the sample selection for the few-shot learning setting. However, running an extensive prompt-engineering would introduce another controllable parameter which could potentially introduce further bias into the results (e.g., different selection strategies showing different impact for different prompt formats).
Finally, we also identify factors that we consider to be sufficiently covered, but where the further extension could lead to more generalisable results (nevertheless, also requiring a considerable amount of additional computational resources), such as including more selection strategies, evaluating on more datasets covering other classification tasks (mainly in regards to text datasets) or using larger number of base models for meta-learning and few-shot fine-tuning, or large language models for in-context learning. First of all, although we focus on the major selection strategies, which are general for any few-shot learning approach, the work can be extended to include larger number of in-context learning specific strategies or strategies that do not perform selection based on scoring of the samples. In addition, even though we run the experiments on 8 image datasets and 6 text datasets, covering different classification tasks with different number of classes, the work could be potentially extended to even larger number of datasets and classification tasks, mainly in case of text modality where we are limited by the number of classes the datasets contain (to sufficiently explore few-shot learning). Finally, although we cover all the main groups of few-shot learning approaches, the work could be extended to cover larger number of meta-learning approaches (where we focus on two of them), base models for meta-learning and few-shot fine-tuning (we focus on one model for image data and one model for text data) and large language models for in-context learning (we focus on two of the most recent and currently best performing large language models Mistral and Zephyr). However, we do not expect significantly different results and findings, only an improvement of their generalisability.
Appendix B Few-Shot Learning Problem Formulation

### 17 Apr 2026 v2

@@ Appendix D Additional Results: Table Results from Text Datasets
Appendix E Additional Experiments: Performance on Image Data
In this appendix, we present the setup and results of the main experiments done on image data. The datasets are chosen similarly to text datasets, where we focus on datasets with different characteristics, such as class imbalance and noisy datasets. Specifically, we use 8 datasets from the Meta-Album (Ullah et al., 2022), specifically LR_AM.DOG and LR_AM.AWA for animal classification, VCL.APL for plane classification, HUM_ACT.ACT_410 for classification of human actions, MNF.TEX_DTD for classification of texture, MCR.PRT and MCR.PNU for classification of microscope images, and PLT.FLW for flower classification.
The results of the main experiments, where we investigate the impact of different single-property sample selection strategies on few-shot learning, as well as provide the comparison with the proposed ACSESS method, are presented in an aggregated form in Table 5 and Figure 5. Similarly to experiments on text data, the results for each individual dataset and model are presented in supplementary material.
| Strategy | Image Data - FSL | ||
| ProtoNet | MAML | FT | |
| Random | |||
| Similarity | |||
| Diversity | |||
| Entropy | |||
| Margin | |||
| Least Confidence | |||
| Loss | |||
| CAL | |||
| CRAIG | |||
| DeepFool | |||
| Forgetting | |||
| Glister | |||
| GraNd | |||
| Herding | |||
| KCenter | |||
| Graph-Cut | |||
| CartographyEasy | |||
| CartographyAmbiguous | |||
| CartographyHard | |||
| CartographyEasy+Ambig. | |||
| ACSESSUniform | |||
| ACSESSWeighted | |||
| ACSESSWith Random | |||
Although the results on the image modality are slightly different, the main findings hold true. Many selection strategies perform worse than the Classic selection (such as hard to learn samples), while only specific ones lead to consistently better performance (e.g., easy to learn samples or those that are forgotten the least).
A specific new finding that we can draw is that the overall benefit of sample selection is higher for noisy datasets and limited for high-quality datasets. Sample selection leads to a considerable performance increase over the Classic selection, on average around percentage points, for the datasets with many noisy samples (as determined by the Cartography) or for datasets with large class imbalance (e.g., difference between lowest and median number of samples per class in few hundred samples), such as HUM_ACT.ACT_410, MCR.PRT or MCR.PNU. At the same time, performance increase is quite limited (on average around percentage points) for the datasets with few noisy samples or no extensive class imbalance, such as the VCL.APL, PLT.FLW or NewsCategory.
When it comes to the proposed ACSESS strategy, we still observe the priority on learnability properties. In the case of the image data, the combination of easy to learn samples (Cartography) that are located on the decision boundary (Margin) provides the most benefit. This represents the combination of the lowest number of strategies we have observed so far, which indicates that the sample selection strategies are well-designed for image data, or that the image datasets contain high-quality samples that are easier to identify.
Finally, when it comes to the impact of increasing the number of shots, or using datasets of different sizes, we still observe similar findings on image data (see Figure 6). When changing the number of shots, the highest performance is achieved around shots and then remains the same across all of further increases. Similarly, when selecting from datasets of smaller sizes, we can observe the same performance when using as few as samples (as compared to using the full dataset). At the same time, we observe that the more sophisticated selection strategies outperform random selection most of the time, while degrading to random selection on a high number of shots.
Appendix F Additional Experiments: Dataset and Model Dependence of Sample Selection Strategies and their Sensitivity to Random Components
@@ Appendix E Additional Experiments: Performance on Image Data
In this appendix, we provide additional insights from ablation studies. We are mainly focus on: 1) how the impact of selection strategies is affected by the used models and datasets, i.e., look at the model and dataset dependence of different selection strategies and configurations of our proposed method (based on dataset specific results); and 2) how sensitive the strategies are to their random components, i.e., look at how the set of selected samples changes when the strategies are run multiple times and how this affects the overall performance (based on the variance from repeated strategy runs in Figure 7).
@@
Specific strategies show large sensitivity to the effects of random seed. Changing the random seed for specific strategies often leads to different set of selected samples and different performance impact. The core set selection strategies are less affected by this randomness. Specific core set strategies (such as Glister or Graph-Cut) consistently select the same set of samples and show no variance in results. Other strategies core set strategies (e.g., Forgetting, DeepFool or Herding) show only small variance. On the other hand, strategies with random initial condition, such as the similarity selection or some of the active learning strategies, are more sensitive to the effects of the random seed. The ACSESS method shows lower sensitivity to repeated runs (with the exception of weighting with random element), but still may lead to different sets of samples based on the random seed (as the standard deviation is higher than in core set strategies, i.e., up to for the ACSESS method as opposed to or for majority of the core set strategies).
@@
Many selection strategies show strong dataset dependence. In many cases, the strategies that perform well on one dataset (such as DeepFool on the HUM_ACT.ACT_410 dataset) may perform quite poorly on other datasets (such as MCR.PRT or MCR.PNU). Due to this dependence, such strategies cannot be used on any dataset without first evaluating them how well they perform on it. Only few strategies perform consistently on the majority of the datasets, such as Herding showing poor performance on all image datasets, or Cartography showing good performance across all datasets and models (while the most beneficial configuration of this strategy is dependent on data modality). Finally, we do not observe a significant dataset dependence of the ACSESS method, as it identifies the well-performing strategies and leverages their strengths for the specific dataset. This observation holds through for the weighting schemes as well (i.e., both uniform and weighted combinations show consistent performance across datasets).
@@
Many selection strategies show strong model dependence and sensitivity to selected samples. The strategies performing well for one model (GraNd on MCR.PNU for MAML, or KCenter on 20 News for Zephyr) may result in a decrease of performance for other models (GraNd for Few-Shot Fine-Tuning, or KCenter for Mistral on the same setting). We do not observe such strong dependence for our proposed method ACSESS, with the exception for Few-Shot Fine-Tuning, where the uniform combination is always better than the weighted combination. Overall, having a different set of selected samples, either as a result from using different strategy or from repeated runs of the same strategy, affects the Few-Shot Fine-Tuning less, often leading to smaller benefit. At the same time, the MAML method and the in-context learning models (Mistral and Zephyr) are more affected by the selected samples, leading to more varied results and often higher performance increase. Such sensitivity to sample selection was observed in previous works as well (Pecher et al., 2024; Köksal et al., 2022; Zhang et al., 2022; Li and Qiu, 2023). Curiously, Prototypical Networks gain quite large performance increase from sample selection, but show low sensitivity to the selected samples and variance from repeated runs of sample selection. As such, sample selection is more important for MAML and In-Context Learning and less important for Few-Shot Fine-Tuning.
Appendix G Additional Experiments: Comparison of Sample Selection Impact Between 5-Shot and 10-Shot Setting
To better explore the effect of increasing number of selected shots on the sample selection strategies, we repeat the evaluation of all the selection strategies in the 10-shot setting (using the same setup and methodology) and compare the results to 5-shot setting. For both 5-shot and 10-shot setting, we provide a distribution of the results as a difference to the Classic selection and illustrate the comparison of these distributions in Figure 8.
Overall, when moving from 5-shot to 10-shots, we observe an increase in the overall performance of the models for all the selection strategies (with the exception of some bad performing ones) as well as the Classic selection. At the same time, the increase in the performance of the Classic selection is higher, leading to lower difference and lower impact of the different evaluated selection strategies. This finding holds for all strategies across all models and few-shot learning approaches.
All in all, these experiments further reinforce the findings from comparison of ACSESS and Random selection, where we observe that the the sample selection progressively leads to lower impact as we increase the number of shots we select.
Appendix H Additional Experiments: Impact of Increasing Number of Shots for Individual Datasets
In this appendix, we provide results from investigating the impact of increasing the number of shots on the benefit of sample selection separately for the individual dataset. For each dataset, we increase the number of selected shots, starting from 1-shot up to the full dataset. The datasets HUM_ACT.ACT_410 and MCR.PRT contain only up to 50 samples per class when all samples are used and the MNF.TEX_DTD only up to 100. All the remaining datasets either contain at maximum 200 samples, or are subsampled to 200 samples per class if necessary (this is especially done for the text datasets, which are often significantly larger). For in-context learning, we increase the number of shots only up to 50 samples per class, as increasing the number further resulted in significantly lower performance (mainly due to limited context size of the large language models). The results for the image datasets are included in Figure 11, for the text datasets using the gradient few-shot learning in Figure 10 and for text datasets using the in-context learning in Figure 9.
For the majority of the datasets, we observe that the highest performance of the sample selection using the ACSESS method is achieved on lower number of shots and then either remains the same until the full dataset, or gradually decreases as we use more samples. Only for specific datasets (LR_AM.DOG and LR_AM.AWA datasets for all approaches; PLT.FLW and SNIPS for MAML; and MNF.TEX_DTD for Prototypical Networks and MAML) we observe continual increase of performance when using sample selection as we increase the number of shots, with the highest performance being achieved only at the full dataset (or slightly beforehand). In addition, on specific datasets, not all models behave the same, for example, on the MNF.TEX_DTD dataset, only the Few-Shot Fine-Tuning approach achieves highest performance on lower number of shots (which then monotonically decrease as we increase the number of used samples), while the MAML and Prototypical Networks meta-learning approaches achieve the highest performance only when using almost the full dataset.
On the other hand, using the Random selection, we often observe monotonic increase in performance as the number of the selected shots increases, with the highest performance being achieved on the full dataset. This behaviour can be observed across all datasets and few-shot learning approaches (with exception of in-context learning). As such, this further reinforces our findings and main motivation that the data-centric approach can be beneficial for the few-shot learning. Curating a subset of the most informative and high-quality samples can often lead to performance on par or better than the one when using all the samples for training.
Finally, we also observe that the benefit of sample selection is more significant at lower number of shots. The difference in performance between the ACSESS method and the Random selection on the lower number of shots is significantly higher and slowly decreases as we use more shots for training. The exception is MAML on the PLT.FLW and SNIPS datasets, or Few-Shot Fine-Tuning on PLT.FLW and MCR.PNU datasets, where the performance of the sample selection is similar to the one using random selection even when using small number of shots. After a certain point, the sample selection using the ACSESS method degrades to Random selection. As such, the focus on selecting high quality samples is paramount when selecting only few labelled samples per class, while random selection can be safely used when using higher number of shots.
Appendix I Additional Experiments: Selection From Datasets of Different Sizes
In this section, we analyse how the benefit of sample selection strategies changes as we decrease the size of the dataset available. For this ablation study, we apply the ACSESS method and Random selection to select 5 samples per class from dataset subsets of different sizes and compare their benefit to the selection from the full dataset. Each dataset is subsampled into multiple sizes, starting with 5 samples per class up to the full dataset. Each subsampling is repeated 10 times and an average result over these repeats are reported. The full-dataset setting represents the selection from 200 samples per class for each datasets (for datasets with a lower number of samples per class, such as HUM_ACT.ACT_410, MCR.PRT or MNF.TEX_DTD, we observe a constant line after the full dataset was achieved). In all cases we perform selection of 5 shots per class using the ACSESS method and Random selection. Results aggregated over all datasets are in Figure 4 and for individual datasets for the image datasets are included in Figure 14, for the text datasets using the gradient few-shot learning in Figure 13 and for text datasets using the in-context learning in Figure 12.
Sample selection is beneficial even when using dataset with low number of samples. Using only 25% (50 samples per class) for in-context learning or 10% of the dataset (20 samples per class) for gradient few-shot learning, sample selection can still identify samples that lead to similar performance as is achieved when using the full dataset. Decreasing the size further, the benefit of sample selection starts to decrease as well. Using only 10 samples per class, the benefit decreases by approximately (e.g., from performance increase of percentage points to only ), and then quickly deteriorates to the Random selection (as the number of samples to choose from is too low).
For majority of the combinations of datasets and approaches, we can observe that sample selection can achieve similar performance using only a fraction of the dataset. For the gradient few-shot learning, this performance is achieved using only 10% of the dataset (20 sample per class), while for in-context learning, this performance is achieved using 25% of the dataset (50 samples per class). However, we observe that this strongly depends on the dataset and the model/approach used. For example on the ATIS dataset and in-context learning, using only 10 samples per class is enough to achieve the same performance as when selecting from the full dataset, while on the SNIPS dataset the number of samples per class increase up to 100. Similarly, on the ATIS dataset with gradient few-shot learning, the MAML a Few-Shot Fine-Tuning approaches achieve the similar performance using only 15-20 samples per class, while Prototypical Networks require the full dataset. Curiously, in specific cases, we can observe that running sample selection on subset of the dataset can lead to better performance than when performing selection from the full dataset (such as on the News Category dataset and gradient few-shot learning).
At the same time, we observe no change in impact of random selection from datasets of different sizes. In this case, the performance remains more or less constant (and often significantly below selection using the ACSESS method) throughout all the sizes – only difference in performance stems from the random subsampling. Finally, using only 10 samples per class, we can observer that sample selection using the ACSESS method can lead to quite big performance improvement over the random selection. Therefore we can conclude that the sample selection is beneficial even when using a small fraction of the dataset size as the one used in the main experiments, although it will lead to a lowered benefit of the sample selection and its impact on the overall performance (which is still higher than when using random selection).
Appendix J Few-Shot Learning Problem Formulation